How many square tiles, each with a perimeter of 12 feet must be used to completely cover a floor with a width of 9 feet and a length of 15 feet?
Answer by stanbon(75887)

If the tiles are square and have a perimeter of 12 feet,each side is 3 ft?
Each tile has area = 3*3 = 9 sq. ft.
-----------
The area of the floor is 9*15 = 135 sq. ft
-----------
The number of tiles needed to cover the floor is 135/9 = 15
=========================
Stan H.